OpenStudy (3psilon):

At t = 0 , a stone is dropped from a cliff above a lake; 1.5 seconds later another stone is thrown downward from the same point with an initial speed of 32 m/s . Both stones hit the water at the same instant. Find the height of the Cliff

OpenStudy (3psilon):

Does the second stone have an acceleration of -g also?

OpenStudy (3psilon):

For the first equation I get \[X _{f} = X _{i}\] is this correct?

OpenStudy (3psilon):

@Shane_B Can you help please?

OpenStudy (3psilon):

@CliffSedge ?

OpenStudy (anonymous):

Did you draw a diagram first?

OpenStudy (3psilon):

Yeah , I'm trying to see what would happen when the first stone is at t = 1.6. Not sure if that'll help though

OpenStudy (anonymous):

No, consider that both are going the same distance, but one is doing it in 1.5s less time.

OpenStudy (anonymous):

You can try setting up two x=vt+0.5t^2 equations where x is the same for both and solve for t.

OpenStudy (anonymous):

oops, x=vt+0.5(a)t^2 [forgot the acceleration]

OpenStudy (3psilon):

So is x = ( x final - x initial )?

OpenStudy (anonymous):

Yes, but x final = 0. (your choice)

OpenStudy (shane_b):

Maybe this will help:\[\frac{1}{2} gt^2=(32m/s)(t-1.5)+\frac{1}{2} gt^2\]\[t=\frac{3}{2}s\]

OpenStudy (anonymous):

Hmm, remember that the second one happened in 1.5s less time.

OpenStudy (anonymous):

What about the second term?

OpenStudy (shane_b):

Yea...brainfart Corrected:\[\frac{1}{2} gt^2=(32m/s)(t-1.5s)+\frac{1}{2} g(t-1.5s)^2\]\[t=2.13728s\]

OpenStudy (3psilon):

But at 1.5 seconds before t = 0 , But why does the equation start with (1/2)gt^2? Wouldn't it equal zero?

OpenStudy (shane_b):

The left side is the stone that is dropped...the right side is the stone that is thrown downward at 32m/s, 1.5s later.

OpenStudy (shane_b):

If they hit at the same time, the distances must be equal. All you need to do now is plug t into either side of the equation to get the distance.

OpenStudy (3psilon):

Ok ok . Sorry but idk if I'm not seeeing it right but the stone dropped was dropped at t = 0 though

OpenStudy (anonymous):

Yes, that's your base line so nothing changes in the basic equation.

OpenStudy (shane_b):

Yes...let's look at it one by one. The stone that was dropped at t=0s. The distance it had to fall would be:\[d=\frac{1}{2}gt^2\]The second stone was thrown 1.5s later...and fell the same distance in basically 1.5s less time. Since it had an initial velocity the distance for that stone would be: \[d=V_0(t-1.5s)+\frac{1}{2}g(t-1.5s)^2\]If they hit the water at the same instant, these distances must be equal so you can set them equal to solve for t...which is what i did above.

OpenStudy (shane_b):

Once you know t you can solve for the distance:\[d=\frac{1}{2}gt^2=\frac{1}{2}(9.8m/s^2)(2.13728s)^2=22.38m\]

OpenStudy (3psilon):

Ok ok I can see it. So the X initial is basically a baseline also?

OpenStudy (shane_b):

I didn't use any formulas with Xi or Xf...just d=V0t+(1/2)gt^2

OpenStudy (shane_b):

I guess you could say d=Xf-Xi...but there's no point :)

OpenStudy (3psilon):

Oh ok I see what you mean! Thank you @Shane_B You've really been a lot of help.

OpenStudy (shane_b):

Np, good luck :)